INAF LBC+LUCIFER queue observing, 2011 June 25 -26

Observer: R. Speziali, K. Boutsia
Telescope Operator: G. Bechetti
Support Astronomer: D. Thompson
Instrument: LUCIFER

Summary

Clear sky at the beginning of the night
The seeing is variable between 0.9''-1.4'' in the optical (DIMM and guider values) - seeing measured on the acquisition images <1.0''

Data Summary

Project Objects Filter ExpTime Notes
Maiolino HDF 210_zJHK - H 2hr arc + flat + Tell
Magrini 7c17_cluster 210_zJHK - H 2hr arc+flat + Tell (?)
Magrini 7c17_cluster 210_zJHK - J 0.5hr arc+flat + Tell (?)

Details

02:11 - Initialize

02:37 - Twilight flats Ks,J,H
  • Ks : 0005-0014 (13000 - 11000cnts) / 0015-0024 (7000-6000cnts) / 0025-0034 (5000-4500cnts)
  • H : 0035-0044 (16000 - 12000cnts) / 0055-0064 (6000-4000cnts) / 0075-0084 (3000-1700cnts)
  • J : 0045-0054 (17000 - 13000cnts) / 0065-0074 (6000-4000cnts) / 0085-0094 (2000-1700cnts)

03:01 - Correct pointing

03:19 - Tell_HDF_H1_acq 0095-0101
  • again we had to shift image manually to center the telluric
  • this is because the N3.75 camera is not well centered and we could not see half of the slit where the telluric was supposed to go
  • some images were saved outside the script
  • DIMM seeing 1.2''

03:39 - Tell_HDF_H1_science 0102-0108

03:46 - HDF_H1_acquisition script 0109-0111
  • sources are too faint - we need to increase exp. time

03:57 - the guide star drifted outside the GCS field of view... need to correct pointing again

04:04 - HDF_H1_acquisition script 0112-0116
  • exptime increased to DIT=20 and NDIT=6 and still 2 of the reference stars were not bright enough

04:26 - science_HDF_H1_1 0117-0128
  • DIMM seeing 1.0'' - guider seeing 0.7''

05:33 - science_HDF_H1_2 0129-0140
  • offset Y=-1.0 to center mask

06:04 - Calib Unit IN
  • ARC 0141-0148
  • FLAT 0149-0154

06:54 - Calib Unit OUT
  • correct pointing

07:00 - Tell_HDF_H1_acquisition (lower airmass) 0155-0158

07:12 - Tell_HDF_H1_science 0159-0165
  • the telluric is centered at the lower half of the slit and in the short slits, part of the light might be lost

07:20 - magrini_acq_sci.txt 0166-0170
  • DIMM 1.4'' - guider 1.4'' - seeing measured on acquisition images 1.0''

07:46 - magrini_sci_H4.txt 0171-0182
  • DIMM 1.3'' - guider 1.1''

08:53 - magrini_sci_H5.txt 0183-0194

09:59 - Calib Unit IN
  • ARC 0195 - 0206
  • FLAT 0207 - 0215

10:10 - Calib Unit OUT

10:11 - magrini_tell_acq 0216-0217
  • manually put the source in the slit
  • the telluric is saturated with 2 sec in Brg !!

10:20 - tell_int_H 0218-0221
  • the spectrum of the telluric seems saturated
  • we change exposure time and repeat
  • also the offsets between the slits are not correct and we can see the telluric only at the initial position

10:29 - tell_acq_ 0222

10:34 - tell_int_H 0223-0226
  • exposure time changed to 3secs

10:36 - tell_acq_again to do filter J 0227

10:41 - tell_int_J 0228-0231
  • also here we changed exposure time to 3sec

10:45 - magrini_spec_acq for filter J 0232-0236
  • we manually apply the offsets calculated yestarday and the mask seems well aligned

11:06 - magrini_spec_sci_J 0237-0242
  • background is getting too high
  • we might not be guiding during the last two exposures

11:42 - Calib Unit IN
  • ARC 0243-0254
  • FLAT 0255-0299

11:57 - Calib Unit OUT

11:58 - End of night - DoDarks script

-- KonstantinaBoutsia - 26 Jun 2011
Topic revision: r6 - 27 Jun 2011, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback